# Efficient Classical Algorithms for Simulating Gaussian Boson Sampling on Graphs

### v1

@@ C.3 Complete Graph $G_{2n}$
  • For complete graph $G_{2n}$, symmetry condition stated in Proposition 16 still holds. We restate it here for clarity:
Proposition 21.
For complete graphs, under the aforementioned multiple canonical paths construction in Appendix C.1, the congestion of each transition in $T^{+}(k)$ (or $T^{-}(k)$) is identical.
Now we estimate the congestion of one transition $t$ in complete graph $G_{2n}$. First, we calculate $\pi(M)$ for each matching. Since any matching $M$ expands a complete subgraph of size $2|M|$, we have $\operatorname{Haf}(M)=(2|M|-1)!!$!, and thus $\pi(M)\propto\lambda^{2|M|}(2|M|-1)!!$!. Then for edge sets of any two matchings $M$ and $M\cup\{e\}$, we have
$\Pr[M\to M\cup\{e\}]$ $=\frac{1}{n(2n-1)}\cdot\frac{\lambda^{2}}{1+\lambda^{2}},$
$\Pr[M\cup\{e\}\to M]$ $=\frac{1}{n(2n-1)}\cdot\frac{\operatorname{Haf}(G_{M})}{\operatorname{Haf}(G_{M\cup\{e\}})}\cdot\frac{1}{1+\lambda^{2}}\geq\frac{1}{n(2n-1)}\cdot\frac{1}{2|M|+1}\cdot\frac{1}{1+\lambda^{2}}.$
@@
$N_{p}=$ $\binom{2n}{2p}\cdot(2p-1)!!=\binom{2n}{2p}\cdot\frac{(2p)!}{2^{p}p!}=\frac{(2n)!}{p!(2n-2p)!\cdot 2^{p}}$
$N_{p,q}=$$\begin{cases}N_{p}\cdot\binom{2n-2p}{2}&\text{ if }q=p+1;\\ N_{p}\cdot p&\text{ if }q=p-1.\end{cases}$
@@
$\quad N_{k,k+1}\cdot\varrho_{t}\leq$ $2\cdot\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{I:|I|\leq k+1}\sum_{F:|F|\geq k-1}\pi(I)\pi(F)|\gamma_{IF}|^{2}$
$\leq$ $2\cdot\frac{n(2n-1)(1+\lambda^{2})}{\lambda^{2}}\cdot(2n)^{2}\cdot N_{k,k+1}\cdot\frac{1}{N_{k,k+1}}\cdot\frac{\sum_{l=1}^{k+1}N_{l}\cdot\lambda^{2l}l!\cdot\sum_{l=k-1}^{n}N_{l}\cdot\lambda^{2l}l!}{\lambda^{2k}k!\cdot\sum_{l=1}^{n}N_{l}\cdot\lambda^{2l}l!}.$
Similarly, notice that $\frac{N_{l+1}\cdot\lambda^{2(l+1)}(l+1)!}{N_{l}\cdot\lambda^{2l}l!}=\frac{(2n-2l)(2n-2l-1)}{2(l+1)}\cdot\lambda^{2}\cdot(l+1)\geq 6\lambda^{2}$ for $l<n-1$, when $\lambda^{2}>\frac{1}{6}$, we have
$\frac{\sum_{l=1}^{k+1}N_{l}\cdot\lambda^{2l}l!}{N_{k,k+1}\lambda^{2k}k!}=$ $\frac{2}{(2n-2k)(2n-2k-1)}\cdot\frac{\sum_{l=1}^{k+1}N_{l}\lambda^{2l}l!}{N_{k+1}\cdot\lambda^{2(k+1)}(k+1)!}\cdot\frac{N_{k+1}\lambda^{2}(k+1)}{N_{k}}$
$\leq$ $\frac{2}{(2n-2k)(2n-2k-1)}\cdot\left(\frac{6\lambda^{2}}{6\lambda^{2}-1}+1\right)\cdot\frac{1}{2}\lambda^{2}(2n-2k)(2n-2k-1)$
$=$ $\lambda^{2}\left(\frac{6\lambda^{2}}{6\lambda^{2}-1}+1\right).$
@@
$N_{k,k-1}\cdot\varrho_{t}\leq$ $2\cdot\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{I:|I|\leq k}\sum_{F:|F|\geq k-1}\pi(I)\pi(F)|\gamma_{IF}|^{2}$
$\leq$ $2\cdot n(2n-1)(2n+1)(1+\lambda^{2})\cdot(2n)^{2}\cdot N_{k,k-1}\cdot\frac{1}{N_{k,k-1}}\cdot\frac{\sum_{l=1}^{k}N_{l}\cdot\lambda^{2l}l!\cdot\sum_{l=k-1}^{n}N_{l}\cdot\lambda^{2l}l!}{\lambda^{2k}k!\cdot\sum_{l=1}^{n}N_{l}\cdot\lambda^{2l}l!}.$
@@
$\frac{\sum_{l=1}^{k}N_{l}\cdot\lambda^{2l}l!}{N_{k,k-1}\lambda^{2k}k!}=$ $\frac{1}{k}\cdot\frac{\sum_{l=1}^{k}N_{l}\cdot\lambda^{2l}l!}{N_{k}\cdot\lambda^{2k}k!}\leq\frac{1}{k}\cdot\left(\frac{6\lambda 62}{6\lambda^{2}-1}+1\right)\leq\frac{6\lambda^{2}}{6\lambda^{2}-1}+1.$
@@
$\varrho_{t}$ $\leq\max\left\{2\cdot\frac{n(2n-1)(1+\lambda^{2})}{\lambda^{2}}\cdot(2n)^{2}\cdot\lambda^{2}\left(\frac{6\lambda^{2}}{6\lambda^{2}-1}+1\right),2\cdot n(2n-1)(2n+1)(1+\lambda^{2})\cdot(2n)^{2}\cdot\left(\frac{6\lambda^{2}}{6\lambda^{2}-1}+1\right)\right\}$
$\leq 32n^{5}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}.$
@@
By Theorem 10, we establish a polynomial mixing time of double-loop Glauber on complete graphs:
Lemma 22.
@@
$\eta_{\frac{\epsilon}{2}}=\frac{\epsilon}{2}/r_{\frac{\epsilon}{2}}=\frac{\epsilon}{2}\left(64n^{5}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}\cdot\left(\ln{\epsilon^{-1}}+n\ln|{\lambda}|+3n\ln{(2n+1)}\right)\right)^{-1}$
in Lemma 5. Then the failure probability of the inner layer sampling in each iteration can be bounded within $\eta_{\frac{\epsilon}{2}}$. Again by Lemma 22, after $r_{\frac{\epsilon}{2}}$ iterations, the total variance can be bounded by $\frac{\epsilon}{2}+\eta_{\frac{\epsilon}{2}}\cdot r_{\frac{\epsilon}{2}}=\epsilon$. Therefore, we give a polynomial-time upper bound for the mixing time of the double-loop Glauber dynamics on the complete graph $G_{2n}$:
Theorem 23.
@@
C.4 Dense Non-bipartite Graphs
In this subsection, we explore extending the results from Appendix C.3 to non-complete graphs, providing proofs of Theorem 3 and Theorem 7. Similar as Appendix C.2, we discuss some dense graphs $G$: there exists a constant $\xi$, such that each vertex is non-adjacent to at most $\xi$ vertices.
For any graph $G$, let $G^{\prime}$ be the corresponding complete bipartite graph with the same vertex set as $G$. Let $\varrho_{t}^{\prime}$ denote the congestion of transition $t=(M,M^{\prime})$ in $G^{\prime}$, $P(M,M^{\prime})$ denote the corresponding transition probability of double-loop Glauber for $G^{\prime}$, and $\operatorname{Haf}^{\prime}(M)$ denote the corresponding Hafnian of the subgraph induced by matching $M$ in $G^{\prime}$, which is equal to $(2|M|-1)!!$!. Then we use the estimate of $\varrho_{t}^{\prime}$ given in Appendix C.3 to bound $\varrho_{t}$ in the original graph $G$.
@@ C.4 Dense Non-bipartite Graphs
Note that the minimum degree $d^{\prime}$ of the subgraph $G_{E}$ satisfies $d^{\prime}\geq 2n^{\prime}-1-\xi$. We start from a vertex $u_{1}$, which has $2n^{\prime}-1-\xi$ possible pairing choices with other vertices. For the choice of pair $(u_{1},v_{1})$, we delete the edge $(u_{1},v_{1})$ and all edges incident to $u_{1}$ and $v_{1}$. The remaining graph has $2n^{\prime}-2-\xi$ vertices, and the minimum degree of this graph is at least $2n^{\prime}-3-\xi$. We can repeat this process and find another pair $(u_{2},v_{2})$, and delete the edge $(u_{2},v_{2})$ and all edges incident to $u_{2}$ and $v_{2}$. This procedure can be repeated until we find $n^{\prime}-\xi-1$ pairs of vertices, and then the remaining graph has $12\xi+$ 1 vertices, with the minimum degree of the current graph being at least $12n^{\prime}-1-\xi-2(n^{\prime}-\xi-1)=\xi+$ 1. Thus again by Ore’s theorem, the remaining graph contains a Hamilton cycle. Finally, we choose this Hamilton cycle and find a perfect matching of $G_{E}$. The number of choices is
$(2n^{\prime}-\xi-1)\cdot(2n^{\prime}-\xi-3)\cdots(\xi+3)=\frac{(2n^{\prime}-1-\xi)!!}{(\xi+1)!!}.$
@@
$1\geq\frac{\operatorname{Haf}(M)}{\operatorname{Haf}^{\prime}(M)}\geq\frac{{(2n^{\prime}-1-\xi)!!}/{(\xi+1)!!}}{(2n^{\prime}-1)!!}\geq\left((2n^{\prime})^{\frac{\xi+1}{2}}\cdot(\xi+1)!!\right)^{-1}\geq\frac{1}{(2n)^{\xi}}.$
@@
ϱ=italic-$\varrho=$ $\max_{t=(M,M^{\prime})}\bigg{\{}\frac{\sum_{(I,F,i)\in cp(t)}p^{(i)}(I,F)\cdot\lambda^{2|I|}\operatorname{Haf}(I)\lambda^{2|F|}\operatorname{Haf}(F)|\gamma_{IF}|}{\lambda^{2|\tilde{M}|}\operatorname{Haf}(\tilde{M})\cdot\sum_{X\text{ is a matching}}\lambda^{2|X|}\operatorname{Haf}(X)}\cdot|E|\frac{1+\lambda^{2}}{{\lambda}^{2}}\bigg{\}}$
$\leq$ $\max_{t=(M,M^{\prime})}\bigg{\{}\frac{\sum_{(I,F,i)\in cp(t)}p^{(i)}(I,F)\cdot\lambda^{2|I|}\operatorname{Haf}^{\prime}(I)\lambda^{2|F|}\operatorname{Haf}^{\prime}(F)|\gamma_{IF}|}{\lambda^{2|\tilde{M}|}\operatorname{Haf}^{\prime}(\tilde{M})(2n)^{-\xi}\cdot\sum_{X\text{ is a matching}}\lambda^{2|X|}\operatorname{Haf}^{\prime}(X)(2n)^{-\xi}}\cdot|E^{\prime}|\frac{1+\lambda^{2}}{{\lambda}^{2}}\bigg{\}}$
$\leq$ ${(2n)}^{2\xi}\varrho^{\prime}.$
@@
By the analysis of Lemma 22, we have the congestion $\varrho^{\prime}$ of double-loop Glauber on complete graph $G_{2n}$ is bounded by $\varrho^{\prime}=32n^{5}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}$. Then we have the congestion $\varrho$ of double-loop Glauber on bipartite graph $G$ is bounded by $\varrho\leq{2n}^{2\xi}\cdot\varrho^{\prime}=32\cdot 2^{\xi}\cdot n^{5+2\xi}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}$. Thus by Theorem 10, the mixing time of double-loop Glauber on complete graph $G_{2n}$ is less than $64\cdot 2^{\xi}\cdot n^{5+2\xi}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}\cdot\left(\ln{\epsilon^{-1}}+n\ln|{\lambda}|+3n\ln{(2n+1)}\right)$. ∎
Finally, we complete the proof of Theorem 7 by extending the analysis of Theorem 23 to the case of dense graphs.
@@
$\eta_{\frac{\epsilon}{2}}=\frac{\epsilon}{2}\left(64\cdot 2^{\xi}\cdot n^{5+2\xi}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}\cdot\left(\ln{\epsilon^{-1}}+n\ln|{\lambda}|+3n\ln{(2n+1)}\right)\right)^{-1}$
in Lemma 4. Then the failure probability of each inner layer sampling step can be bounded within $\eta_{\frac{\epsilon}{2}}$, and need $\tilde{O}(n^{14}(\log{\epsilon^{-1}})^{2})$ time. Again by Lemma 22, after
$64\cdot 2^{\xi}\cdot n^{5+2\xi}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}\cdot\left(\ln{\epsilon^{-1}}+n\ln|{\lambda}|+3n\ln{(2n+1)}\right)$
@@
Appendix D Double-loop Glauber Dynamics on Weighted Graphs

### v2

@@ C Complete Graph $G_{2n}$
  • For complete graph $G_{2n}$, symmetry condition stated in Proposition S1 still holds. We restate it here for clarity:
Proposition S3.
For complete graphs, under the aforementioned multiple canonical paths construction in Supplementary Note A, the congestion of each transition in $T^{+}(k)$ (or $T^{-}(k)$) is identical.
Now we estimate the congestion of one transition $t$ in complete graph $G_{2n}$. First, we calculate $\pi(M)$ for each matching. Since any matching $M$ expands a complete subgraph of size $2|M|$, we have $\operatorname{Haf}(M)=(2|M|-1)!!$, and thus $\pi(M)\propto\lambda^{2|M|}(2|M|-1)!!$. Then for edge sets of any two matchings $M$ and $M\cup\{e\}$, we have
$\Pr[M\to M\cup\{e\}]$ $=\frac{1}{n(2n-1)}\cdot\frac{\lambda^{2}}{1+\lambda^{2}},$ (S60)
$\Pr[M\cup\{e\}\to M]$ $=\frac{1}{n(2n-1)}\cdot\frac{\operatorname{Haf}(G_{M})}{\operatorname{Haf}(G_{M\cup\{e\}})}\cdot\frac{1}{1+\lambda^{2}}\geq\frac{1}{n(2n-1)}\cdot\frac{1}{2|M|+1}\cdot\frac{1}{1+\lambda^{2}}.$ (S61)
@@
$N_{p}=$ $\binom{2n}{2p}\cdot(2p-1)!!=\binom{2n}{2p}\cdot\frac{(2p)!}{2^{p}p!}=\frac{(2n)!}{p!(2n-2p)!\cdot 2^{p}}$ (S62)
[…]
@@
$\quad N_{k,k+1}\cdot\varrho_{t}\leq$ $2\cdot\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{I:|I|\leq k+1}\sum_{F:|F|\geq k-1}\pi(I)\pi(F)|\gamma_{IF}^{(i)}|^{2}$ (S64)
$\leq$ $2\cdot\frac{n(2n-1)(1+\lambda^{2})}{\lambda^{2}}\cdot(2n)^{2}\cdot N_{k,k+1}\cdot\frac{1}{N_{k,k+1}}\cdot\frac{\sum_{l=1}^{k+1}N_{l}\cdot\lambda^{2l}l!\cdot\sum_{l=k-1}^{n}N_{l}\cdot\lambda^{2l}l!}{\lambda^{2k}k!\cdot\sum_{l=1}^{n}N_{l}\cdot\lambda^{2l}l!}.$ (S65)
Similarly, notice that $\frac{N_{l+1}\cdot\lambda^{2(l+1)}(l+1)!}{N_{l}\cdot\lambda^{2l}l!}=\frac{(2n-2l)(2n-2l-1)}{2(l+1)}\cdot\lambda^{2}\cdot(l+1)\geq 6\lambda^{2}$ for l<n1l<n-1, when $\lambda^{2}>\frac{1}{6}$, we have
$\frac{\sum_{l=1}^{k+1}N_{l}\cdot\lambda^{2l}l!}{N_{k,k+1}\lambda^{2k}k!}=$ $\frac{2}{(2n-2k)(2n-2k-1)}\cdot\frac{\sum_{l=1}^{k+1}N_{l}\lambda^{2l}l!}{N_{k+1}\cdot\lambda^{2(k+1)}(k+1)!}\cdot\frac{N_{k+1}\lambda^{2}(k+1)}{N_{k}}$ (S66)
$\leq$ $\frac{2}{(2n-2k)(2n-2k-1)}\cdot\left(\frac{6\lambda^{2}}{6\lambda^{2}-1}+1\right)\cdot\frac{1}{2}\lambda^{2}(2n-2k)(2n-2k-1)$ (S67)
$=$ $\lambda^{2}\left(\frac{6\lambda^{2}}{6\lambda^{2}-1}+1\right).$ (S68)
@@
$N_{k,k-1}\cdot\varrho_{t}\leq$ $2\cdot\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{I:|I|\leq k}\sum_{F:|F|\geq k-1}\pi(I)\pi(F)|\gamma_{IF}^{(i)}|^{2}$ (S69)
$\leq$ $2\cdot n(2n-1)(2n+1)(1+\lambda^{2})\cdot(2n)^{2}\cdot N_{k,k-1}\cdot\frac{1}{N_{k,k-1}}\cdot\frac{\sum_{l=1}^{k}N_{l}\cdot\lambda^{2l}l!\cdot\sum_{l=k-1}^{n}N_{l}\cdot\lambda^{2l}l!}{\lambda^{2k}k!\cdot\sum_{l=1}^{n}N_{l}\cdot\lambda^{2l}l!}.$ (S70)
@@
$\frac{\sum_{l=1}^{k}N_{l}\cdot\lambda^{2l}l!}{N_{k,k-1}\lambda^{2k}k!}=$ $\frac{1}{k}\cdot\frac{\sum_{l=1}^{k}N_{l}\cdot\lambda^{2l}l!}{N_{k}\cdot\lambda^{2k}k!}\leq\frac{1}{k}\cdot\left(\frac{6\lambda 62}{6\lambda^{2}-1}+1\right)\leq\frac{6\lambda^{2}}{6\lambda^{2}-1}+1.$ (S71)
@@
$\varrho_{t}$ $\leq\max\left\{2\frac{n(2n-1)(1+\lambda^{2})}{\lambda^{2}}(2n)^{2}\lambda^{2}\left(\frac{6\lambda^{2}}{6\lambda^{2}-1}+1\right),2n(2n-1)(2n+1)(1+\lambda^{2})(2n)^{2}\left(\frac{6\lambda^{2}}{6\lambda^{2}-1}+1\right)\right\}$ (S72)
$\leq 32n^{5}\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}.$ (S73)
@@
By Theorem S2, we establish a polynomial mixing time of double-loop Glauber on complete graphs:
Lemma S8.
@@
$\eta_{\frac{\epsilon}{2}}=\frac{\epsilon}{2}/r_{\frac{\epsilon}{2}}=\frac{\epsilon}{2}\left(64n^{5}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}\cdot\left(\ln{\epsilon^{-1}}+n\ln|{\lambda}|+3n\ln{(2n+1)}\right)\right)^{-1}$ (S74)
in Lemma 2. Then the failure probability of the inner layer sampling in each iteration can be bounded within $\eta_{\frac{\epsilon}{2}}$. Again by Lemma S8, after $r_{\frac{\epsilon}{2}}$ iterations, the total variance can be bounded by $\frac{\epsilon}{2}+\eta_{\frac{\epsilon}{2}}\cdot r_{\frac{\epsilon}{2}}=\epsilon$. Therefore, we give a polynomial-time upper bound for the mixing time of the double-loop Glauber dynamics on the complete graph $G_{2n}$:
Theorem S6.
@@
D Dense Non-bipartite Graphs
In this subsection, we explore extending the results from Supplementary Note C to non-complete graphs, providing proofs of Theorem 3 and Theorem 5. Similar as Supplementary Note B, we discuss some dense graphs $G$: there exists a constant $\xi$, such that each vertex is non-adjacent to at most $\xi$ vertices.
For any graph $G$, let $G^{\prime}$ be the corresponding complete bipartite graph with the same vertex set as $G$. Let $\varrho_{t}^{\prime}$ denote the congestion of transition $t=(M,M^{\prime})$ in $G^{\prime}$, $P(M,M^{\prime})$ denote the corresponding transition probability of double-loop Glauber for $G^{\prime}$, and $\operatorname{Haf}^{\prime}(M)$ denote the corresponding Hafnian of the subgraph induced by matching $M$ in $G^{\prime}$, which is equal to (2|M|1)!!(2|M|-1)!!. Then we use the estimate of $\varrho_{t}^{\prime}$ given in Supplementary Note C to bound $\varrho_{t}$ in the original graph $G$.
@@ D Dense Non-bipartite Graphs
Note that the minimum degree $d^{\prime}$ of the subgraph $G_{E}$ satisfies $d^{\prime}\geq 2n^{\prime}-1-\xi$. We start from a vertex $u_{1}$, which has $2n^{\prime}-1-\xi$ possible pairing choices with other vertices. For the choice of pair $(u_{1},v_{1})$, we delete the edge $(u_{1},v_{1})$ and all edges incident to $u_{1}$ and $v_{1}$. The remaining graph has $2n^{\prime}-2-\xi$ vertices, and the minimum degree of this graph is at least $2n^{\prime}-3-\xi$. We can repeat this process and find another pair $(u_{2},v_{2})$, and delete the edge $(u_{2},v_{2})$ and all edges incident to $u_{2}$ and $v_{2}$. This procedure can be repeated until we find $n^{\prime}-\xi-1$ pairs of vertices, and then the remaining graph has $2\xi+1$ vertices, with the minimum degree of the current graph being at least $2n^{\prime}-1-\xi-2(n^{\prime}-\xi-1)=\xi+1$. Thus again by Ore’s theorem, the remaining graph contains a Hamilton cycle. Finally, we choose this Hamilton cycle and find a perfect matching of $G_{E}$. The number of choices is
$(2n^{\prime}-\xi-1)\cdot(2n^{\prime}-\xi-3)\cdots(\xi+3)=\frac{(2n^{\prime}-1-\xi)!!}{(\xi+1)!!}.$ (S76)
@@
$1\geq\frac{\operatorname{Haf}(M)}{\operatorname{Haf}^{\prime}(M)}\geq\frac{{(2n^{\prime}-1-\xi)!!}/{(\xi+1)!!}}{(2n^{\prime}-1)!!}\geq\left((2n^{\prime})^{\frac{\xi+1}{2}}\cdot(\xi+1)!!\right)^{-1}\geq\frac{1}{(2n)^{\xi}}.$ (S77)
@@
$\varrho=$ $\max_{t=(M,M^{\prime})}\bigg\{\frac{\sum_{(I,F,i)\in cp(t)}p^{(i)}(I,F)\cdot\lambda^{2|I|}\operatorname{Haf}(I)\lambda^{2|F|}\operatorname{Haf}(F)|\gamma_{IF}^{(i)}|}{\lambda^{2|\tilde{M}|}\operatorname{Haf}(\tilde{M})\cdot\sum_{X\text{ is a matching}}\lambda^{2|X|}\operatorname{Haf}(X)}\cdot|E|\frac{1+\lambda^{2}}{{\lambda}^{2}}\bigg\}$ (S78)
$\leq$ $\max_{t=(M,M^{\prime})}\bigg\{\frac{\sum_{(I,F,i)\in cp(t)}p^{(i)}(I,F)\cdot\lambda^{2|I|}\operatorname{Haf}^{\prime}(I)\lambda^{2|F|}\operatorname{Haf}^{\prime}(F)|\gamma_{IF}^{(i)}|}{\lambda^{2|\tilde{M}|}\operatorname{Haf}^{\prime}(\tilde{M})(2n)^{-\xi}\cdot\sum_{X\text{ is a matching}}\lambda^{2|X|}\operatorname{Haf}^{\prime}(X)(2n)^{-\xi}}\cdot|E^{\prime}|\frac{1+\lambda^{2}}{{\lambda}^{2}}\bigg\}$ (S79)
$\leq$ ${(2n)}^{2\xi}\varrho^{\prime}.$ (S80)
@@
By the analysis of Lemma S8, we have the congestion $\varrho^{\prime}$ of double-loop Glauber on complete graph $G_{2n}$ is bounded by $\varrho^{\prime}=32n^{5}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}$. Then we have the congestion $\varrho$ of double-loop Glauber on bipartite graph $G$ is bounded by $\varrho\leq{2n}^{2\xi}\cdot\varrho^{\prime}=32\cdot 2^{\xi}\cdot n^{5+2\xi}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}$. Thus by Theorem S2, the mixing time of double-loop Glauber on complete graph $G_{2n}$ is less than $64\cdot 2^{\xi}\cdot n^{5+2\xi}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}\cdot\left(\ln{\epsilon^{-1}}+n\ln|{\lambda}|+3n\ln{(2n+1)}\right)$. ∎
Finally, we complete the proof of Theorem 5 by extending the analysis of Theorem S6 to the case of dense graphs.
@@
$\eta_{\frac{\epsilon}{2}}=\frac{\epsilon}{2}\left(64\cdot 2^{\xi}\cdot n^{5+2\xi}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}\cdot\left(\ln{\epsilon^{-1}}+n\ln|{\lambda}|+3n\ln{(2n+1)}\right)\right)^{-1}$ (S81)
in Lemma 1. Then the failure probability of each inner layer sampling step can be bounded within $\eta_{\frac{\epsilon}{2}}$, and need $\tilde{O}(n^{14}(\log{\epsilon^{-1}})^{2})$ time. Again by Lemma S8, after
$64\cdot 2^{\xi}\cdot n^{5+2\xi}\cdot\frac{(\lambda^{2}+1)(12\lambda^{2}-1)}{6\lambda^{2}-1}\cdot\left(\ln{\epsilon^{-1}}+n\ln|{\lambda}|+3n\ln{(2n+1)}\right)$ (S82)
@@
E Technical Barrier for Non-dense Graphs
In this subsection, we highlight several technical challenges that arise when extending our approach to non-dense graphs.
For the proof for dense graphs in Supplementary Note B and Supplementary Note D, the key insight of our proof relies on bounding the Hafnian of matchings in the given dense graph via the Hafnian of matchings in corresponding complete graphs, as demonstrated in (S54) and (S77). For non-dense graphs, unfortunately the ratio between the Hafnian in the current graph and that in complete graphs can become exponentially large, particularly for sparse graphs. Consequently, directly applying results from complete graphs fails to yield polynomial mixing time.
This observation exposes a fundamental limitation of the canonical path method in our setting. The Hafnian intrinsically corresponds to the weight of matchings in the stationary distribution. While for dense graphs we effectively bound these weights using Hafnian in corresponding complete graphs, for non-dense graphs, even matchings with a large number of edges can have very small Hafnian. Consequently, such low-weight matchings are likely to become the bottleneck in the canonical path construction, inducing large congestion.
Here we consider a potential hard instance shown in Supplementary Fig. 2a. The graph consists of several squares arranged in a cycle. Label the four vertices of each square clockwise, and connect the first vertex of each square to its third vertex with an edge, and connect the second vertex of each square to the fourth vertex of the previous square with an edge.
Supplementary Fig. 2: A hard instance for non-dense graphs and its two different types of perfect matchings. The graph in the left figure consists of several squares arranged in a cycle, where the first vertex of each square is connected to its third vertex with an edge, and the second vertex of each square is connected to the fourth vertex of the previous square by an edge. The red edges in the right figure form a perfect matching $M_{0}$ in set AA (which is also the unique one); the blue edges form a perfect matching $M_{1}$ in set $B$. Notice that deleting any edge in $M_{0}$ will result in a near perfect matching with Hafnian 11.
Notice that the perfect matchings of this graph can be divided into two categories: those that include the edge connecting different squares (denoted as set $A$), and those that only include edges within each square (denoted as set $B$). The first category of perfect matchings has only one perfect matching $M_{0}$, as is shown in Supplementary Fig. 2b. The second category of perfect matchings has $2^{n}$ perfect matchings, where $n$ is the number of squares in the graph. This is because each square has two choices for selecting edges, and there are $n$ squares in total. Therefore the Hafnian of any perfect matching is $1+2^{n}$. Now consider the canonical paths from $B$ to $A$. The final step of these canonical paths must be adding one edge from a near perfect matching and get $M_{0}$, which have $2n$ choices. A significant insight is that these $2n$ near perfect matchings are all very low-weight matchings, with Hafnian being $1$. Then we can consider the sum of congestion of these transitions:
$\sum_{e\in M_{0}}\rho_{t=(M_{0}\backslash\{e\},M_{0})}\geq$ $\sum_{e\in M_{0}}\frac{1}{\pi(M_{0}\backslash\{e\})P(M_{0}\backslash\{e\},M_{0})}\cdot\sum_{\begin{subarray}{c}(I,F,i)\in cp(t)\\ I\in B,F\in A\end{subarray}}p^{(i)}(I,F)\cdot\pi(I)\pi(F)|\gamma_{IF}^{(i)}|$ (S83)
$\geq$ $\frac{1}{\pi(M_{0}\backslash{\{e_{0}\}})\cdot P(M_{0}\backslash\{e_{0}\},M_{0})}\cdot\sum_{I\in B}\pi(I)\pi(M_{0})\cdot n$ (S84)
[…]
Furthermore, we can directly give a lower bound of the mixing time of this hard instance by conductance.
Lemma S10 (Theorem 7.4, [19]).
For a given ergodic, aperiodic, and irreducible Markov chain, let $\Phi$ denote the conductance defined in (S16), we have
$t_{\operatorname{mix}}=t_{\operatorname{mix}}\left(\frac{1}{4}\right)\geq\frac{1}{4\Phi}.$ (S86)
The conductance of the set $S$ is defined to be
$\Phi(S):=\frac{\sum_{i\in S,j\not\in S}p_{ij}\pi_{i}}{\sum_{i\in S}\pi_{i}},$ (S87)
and we have
$\Phi=\min_{S:0<\sum_{i\in S}\pi_{i}\leq\frac{1}{2}}\Phi(S).$ (S88)
Therefore, to establish a lower bound on the mixing time, we only need to find a set $S$ such that its conductance is quite small. Just take $S=A$.
[…]
$=$ $2n\cdot\frac{1}{6n}\cdot\frac{1}{1+2^{n}}\cdot\frac{1}{1+\lambda^{2}}.$ (S92)
This means that for the given hard instance, the double-loop Markov chain satisfies
$t_{\operatorname{mix}}\geq\frac{1}{4\Phi}\geq\frac{1}{4\Phi(A)}=3(1+\lambda^{2})\cdot(1+2^{n}),$ (S93)
which is exponentially large.
For non-dense non-bipartite graphs, there is an another additional technical challenge: the task of uniformly sampling perfect matchings from subgraphs in the inner loop is quite challenging for general non-bipartite graphs.
Supplementary Note 4 Double-loop Glauber Dynamics on Weighted Graphs